OpenStudy (anonymous):

if y =-5 when x =15 find y when x=-1

OpenStudy (ganpat):

y = -5 at x = 15, so rewrite it in following expression.. -5 ----> 15 y -----> -1 y = (-5 * -1 )/ 15 = 5/15 = 1/3... i guess.... @waterineyes , @Hero, @ParthKohli , what say guys ??

OpenStudy (anonymous):

@ganpat it gave me choices like y=-1/5x;1/5 y=-1/3x;2/5 y=1/3x;-1/3 y=-1/3x;1/3

OpenStudy (ganpat):

i guess, C shd be answer..

OpenStudy (anonymous):

thanks